OpenStudy (anonymous):

What is 66 2/3% of 300

zepdrix (zepdrix):

Hey Dani :) Welcome to OpenStudy! \(\Large\rm 66\frac{2}{3}\text{% of }100\text{ is }66\frac{2}{3}\)

zepdrix (zepdrix):

But we have three hundreds. So we'll end up with: \(\Large\rm 66\frac{2}{3}+66\frac{2}{3}+66\frac{2}{3}\)

zepdrix (zepdrix):

Hopefully you understand how to add the fractional parts :)